OpenStudy (anonymous):

What property is shown? –8 + (–15) = (–15) + (–8)

OpenStudy (nikato):

communitive because a+b=b+a

OpenStudy (mathstudent55):

Commutative: a + b = b + a (change of order) Associative: a + (b + c) = (a + b) + c (change of grouping)